OpenStudy (anonymous):

A bat and ball cost a dollar and ten cents. The bat costs a dollar more than the ball. How much does the ball cost? If you know the answer type it! lets see who gets it right!

OpenStudy (anonymous):

ball =x, bat = x +1, therefore x+(x+1) = $1.10

OpenStudy (anonymous):

now find x

OpenStudy (phi):

following up on ace: cost of a ball is x: x+x+1= 1.10 2x+1= 1.10 2x= 0.10 x= 0.05 cost of a ball is 0.05 (1 nickel) the cost of a bat is 1+x= 1 +0.05= 1.05 cost of a ball is $0.05 cost of a bat is $1.05 cost of a bat and a ball is $1.10 I would say ace is correct, and left it to you to show the details.

OpenStudy (anonymous):

yes you are right!

OpenStudy (phi):

algebra rules!